OpenStudy (anonymous):

The function y=f(x) is given. Describe how the graphs of the function can be obtained from the graph of y=f(x). The function is: y+7 = f(x)

OpenStudy (anonymous):

The answer is: "translate the graph of y= f(x) downward 7 units" why so?

OpenStudy (anonymous):

would the equation look like: y = f(x) - 7?

OpenStudy (anonymous):

if you don't mind me saying so this is really a bad question \[y+7\] is not a function, but nvm put \[y+7=f(x)\] so \[y=f(x)-7\] translates down 7

OpenStudy (anonymous):

ohh, thank you